OpenStudy (anonymous):

Which is the equation of the line with slope 5 that contains point (-2, -3) ?

OpenStudy (anonymous):

\[\huge y-y _{1}=m(x-x _{1})\] Use this and transform it into y=mx+b form, where m is the slope, and x1 and y1 are the coordinates.

OpenStudy (anonymous):

i dont understand :/

OpenStudy (anonymous):

y=5x-7

OpenStudy (anonymous):

these are my answer choices : A.) y + 3 = 5(x + 2) B.) y – 2 = 5(x – 3) C.) y – 3 = 5(x – 2) D.) y + 2 = 5(x – 3)

OpenStudy (anonymous):

Just use the formula I told you...(x1,y1) are the co-ordinates your given and m is the slope...

OpenStudy (anonymous):

answer choice a

OpenStudy (anonymous):

Don't just give him the answer -.-

OpenStudy (anonymous):

Otherwise the person is not learning.